OpenStudy (anonymous):

Solve for t: 1/4t=16.

OpenStudy (abdullah1995):

1/(16*4) = t

OpenStudy (abdullah1995):

lol

OpenStudy (anonymous):

@tjdaking is the problem: \[\frac{1}{4}t =15\] or \[\frac{1}{4t}=15\]